Tyrion:

Solve the equation the square root of the quantity x plus 4 minus 3 equals 1 for the variable. Show each step of your solution process.

Tyrion:

yuh

Tyrion:

whikh means wat

jhonyy9:

ATTENTION !!! the square root of the quantity x plus 4 minus 3 equals 1

Tyrion:

huh

jhonyy9:

@surjithayer more attention please

surjithayer:

sorry ,i read something else

jhonyy9:

there is minus 3 not plus 3

surjithayer:

\[\sqrt{x+4}-3=1\] \[\sqrt{x+4}=4\] squaring x+4=16 x=?

jhonyy9:

in this way is more easy and you get real solution

Tyrion:

x=12